OpenStudy (anonymous):

Max earns $1,800 per month, plus a commision of 3% of his sales.He wan ts to earn at least $3,000 this month.What amount of sales does max need this month?

OpenStudy (anonymous):

$1,800+0.03X>=$3,000 0.03X>=$3,000-$1,800 X>=$1200/0.03 X>=$40,000 Atleast $40,000 a month

OpenStudy (anonymous):

hey i got an answer for your question the answer would be 40,000. first, we assume the amount of sales be x. 3% of x + 1800 = total (3000) (3/100)x + 1800 = 3000 solve for x. you'll get x = 40000